OpenStudy (anonymous):

Forty is 25% of what number?

OpenStudy (anonymous):

160

OpenStudy (confluxepic):

@Wippin2015k \(\huge\color{blue}{Welcome~To~OpenStudy!}\) Direct answers are not allowed.

OpenStudy (anonymous):

oh well

OpenStudy (anonymous):

this is not my first time on openstudy smart one

OpenStudy (anonymous):

160 is right :)

OpenStudy (anonymous):

To find the solution to this, you can use this formula \(\rm \dfrac{part}{whole} = \dfrac{percent}{100}\) From looking at this, we know that our part is 40, our whole is unknown, and our percent is 25, so plug in the values \(\rm \dfrac{40}{whole} = \dfrac{25}{100}\) We have to cross multiply 40 by 100 to get 4000. Now we take this 4000 and divide it by 25 to get 160, our answer, so \(\rm \dfrac{40}{160} = \dfrac{25}{100}\)